OpenStudy (abbie):

plans for a stadium are drawn on a coordinate grid. one wall lies on the line y=2x+3. a perpendicular wall passes through the point (4, -7). write the equation of the line that contains the new wall

OpenStudy (solomonzelman):

just gave you that link...

OpenStudy (abbie):

y=-1/2x-5 y=2x-15 y=1/2x-9 y=-2x+1

OpenStudy (solomonzelman):

how do you find a perpendicular slope?

OpenStudy (abbie):

idk

OpenStudy (abbie):

im watching a khan academy video though maybe it could help

OpenStudy (solomonzelman):

I can help explain this right now in a short simple phrase.

OpenStudy (abbie):

go on then

OpenStudy (solomonzelman):

multiplicative inverse of the slope then times '-1' is your perpendicular slope.

OpenStudy (solomonzelman):

@abbie, see? So what is the slope going to be?

OpenStudy (solomonzelman):

for a slope a it is -1/a

OpenStudy (abbie):

ok

OpenStudy (solomonzelman):

so if for A the slope is -1/A, for 2 the slope is _

OpenStudy (abbie):

um 1/2?

OpenStudy (solomonzelman):

close, (it is multiplicative inverse, and then times -1)

OpenStudy (abbie):

-2?

OpenStudy (solomonzelman):

you got the 1/2 part right, but think one more time....

OpenStudy (abbie):

-1/2

OpenStudy (solomonzelman):

Yup! Bingo!

OpenStudy (abbie):

lol ok

OpenStudy (solomonzelman):

K, so lets graph the line y=2x+3. Plug in -1 & 1 for x and connect the points that you get with a strait line.

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

now graph the same but add the point (4, -7) to the graph. (don't connect it to the line yet, just do the same graph as mine and draw this additional point.)

OpenStudy (abbie):

Im having a headache right now I think Ill do it later

OpenStudy (solomonzelman):

 I graphed this on a paper.